a) what volume of water has the same mass as 100 mm3 of gold? What is the specific weight of gold?

b) A submarine is submerged at a depth of 200m and then levels off. How it is achieved that inside the submarine atmospheric pressure is mantained? What are the pressure and total force applied to a hatch 1m wide and 1.5m long?



1
Expert's answer
2021-09-15T11:38:53-0400

a)


"V_w=\frac{19300}{1000}100= 1930\ mm^3\\s_w=19.3"

b)


"p=(1000)(9.8)(200)=2.0\cdot10^6\ Pa\\F=(2.0\cdot10^6)(1)(1.5)=3.0\cdot10^6\ N"